OpenStudy (anonymous):

Solve this linear equation: 5(e-4) = -e/3 Please show how to solve step by step.

OpenStudy (anonymous):

5e - 20 = -e/3 15e - 60 = -e 16e = 60 e = 60/16 = 15/4

OpenStudy (anonymous):

For the second step I multiplied both sides by 3 to get rid of the fraction on the right hand side.